### v5

@@ 2.2 Formulation
For the sake of simplicity, the integrated vehicle and pollster routing model (IVPRM) is stated in the following blocks. Table 1 summarizes all the relevant parameters in the model.
@@
The following sets of binary variables for arc selection are used in the model. First, variable $x_{i,j}^{e,s}$ is equal to one if and only if the arc $(i,j)\in A_{S}\cup A_{W}$ is chosen to be part of some walking path of the pollster $e\in E$ on day $s\in S$. Second, $y_{i,j}^{k,s}$ is one whenever arc $(i,j)\in A_{V}$ is part of the route of vehicle $k\in K$ on day $s\in S$. Third, $z_{i,j}^{e,s}$ indicates if pollster $e\in E$ is transported by some vehicle over arc $(i,j)\in A_{V}$ on day $s\in S$. Fourth, $b_{i}^{e,s}$ (correspondingly $f_{i}^{e,s}$) indicates if a pollster $e\in E$ starts (finishes) a walking path at node $i\in C_{-}$ ($i\in C_{+}$) on day $s\in S$. Fifth, $w_{i}^{e,s}$ indicates if a pollster $e\in E$ takes a break at node $i\in C_{-}$ on day $s\in S$. At last, $u_{s}$ determines if any workload is assigned to pollsters and vehicles on day $s\in S$ for the final aggregated daily plan. Additionally, for each node $i\in C_{-}$, let us define $B_{i}$ and $B_{i+n}$ as the arrival and departure times of a pollster at nodes $i$ and $i+n$, respectively.
@@
The objective function aims to minimize the number of operational days and the number of vehicles and pollsters hired. The first component represents the fixed operational daily cost; meanwhile, the other two components represent the variable costs depending on the number of pollsters and vehicles hired for an operational day. Even though this definition of an objective function is not frequent in Routing Problems, this is justified because the distance walked by a pollster or the travel time spent by a vehicle are irrelevant quantities for this particular application (fixed daily costs of hiring vehicles and pollsters).
@@
The following set of constraints describe routes for each pollster $e\in E$ within each day $s\in S$ in the time horizon.
@@
Constraints (2a) ensure that each store $i\in C_{-}$ is visited exactly once by a pollster on a given day. Constraints (2b) and (2c) identify the edges of walking paths, i.e., if $b_{i}^{e,s}=1$ (resp. $f_{i}^{e,s}=1$), then a walking path must begin (resp. end) at node $i\in C_{-}$ (resp. $C_{+}$). Moreover, constraints (2d) – (2g) guarantee that if node $i$ is the beginning or the end of a walking path, then one vehicle must pick up or deliver exactly one pollster on that node. Finally, constraints (2h) enforce that each pollster leaves the depot at most once in every active day.
@@
The following constraints characterize vehicular routes for pollster transportation.
@@
Constraints (3a) and (3b) ensure that vehicles arrive only to pick up nodes and depart just from delivery nodes. (3c) and (3d) are flow conservation constraints. Furthermore, (3e) determine that each vehicle leaves the depot at most once in every active day. Finally, (3f) establish the linkage between pollster and vehicular routes.
Time management and shift-length
The set of constraints below are introduced to describe time bounds for the duration of a pollster shift and times required to travel and visit each store.
@@
Constraints (4a) require that the departure time from a store $i$ to be at least as large as the arrival time plus the service time and, eventually, the time needed if a pause is taken at that store. Observe that the sum in the right-most term in these constraints equals one if a pause is taken at node $i$ and zero otherwise. Constraints (4b) and (4c) ensure that the arrival time at a store within a route must include the pedestrian or vehicular travel time from the previously visited store, respectively. If node $i$ comes right after the depot in any route, then constraints (4d) provide a lower bound on the arrival time of vehicles and pollsters. Finally, constraints (4e) link variables $B_{i}$ and $u^{s}$ to bound the latest arrival time of any vehicle to the depot for each day $s$. Here, $M$ is a sufficiently large number such that $M>\max\{t,\tau\}+B_{\max}$.
@@
The subsequent constraints ensure that each pollster on duty must take a break starting within the prescribed time frame on a given day.
@@
Constraints (5a) establish that if a break is taken by a pollster at node $i$, then it must occur within the time interval $[T_{0},T_{1}]$. Constraints (5b) guarantee that a pollster takes a break only at visited stores, and constraints (5c) enforce that each pollster on duty takes a break once per day.
Table 1: All the parameters of the proposed model.
Symbol Parameter
$n$ Number of stores
$K$ Set of vehicles
$E$ Set of pollsters
$S$ Set of days in the time horizon
$B_{\max}$ Day length
$P$ Length of pause for pollsters
$Q$ Vehicle capacity
$G:=(V,A)$ Multi-graph
$V:=C\cup\{0,2n+1\}$ Node set of $G$
$C:=C_{-}\cup C_{+}$ Set of nodes associated to stores
$C_{-}:=\{1,\dots,n\}$ Set of copies of the nodes (start of service)
$C_{+}:=\{n+1,\dots,2n\}$ Set of copies of the nodes (end of service)
$A:=A_{S}\cup A_{W}\cup A_{V}$ Arc set of $G$
$A_{S}$ Service arcs connecting each store with its duplicate
$A_{W}$ Walking arcs
$A_{V}$ Vehicle arcs
$t_{a}$ Time spent by a pollster collecting data at a store, for $a\in A_{S}$
$t_{a}$ Pedestrian travel time between stores, for $a\in A_{W}$
$\tau_{a}$ Vehicular travel time between stores, for $a\in A_{V}$
$\kappa_{0}$ Fixed working day operational cost
$\kappa_{1}$ Fixed vehicle hiring cost
$\kappa_{2}$ Fixed pollster hiring cost
$[T_{0},T_{1}]$ Time window in which pollsters must start a break
$M$ Sufficiently large number
@@ 2.3 Symmetry breaking inequalities
In any feasible aggregated daily plan, there is a fixed vehicle servicing a vehicular route, yet the route can be served by any other vehicle in $K$, yielding the same objective value. Similarly, the workload of any pair of pollsters in $E$ can be permuted throughout a daily plan. Moreover, the same happens with the set of days $S$ concerning the feasible daily plans. Then, a set of equivalent solutions might appear by relabeling assigned vehicles and pollsters. The following set of constraints tries to avoid these symmetries. It can be seen as an application of the well-known technique of symmetry elimination [39, 40, 41].
@@
Constraints (6a) – (6c) ensure an ordered assignment of pollsters starting with the one labeled as 00 for each day $s\in S$. Similarly, constraints (6d) – (6f) guarantee an ordered assignment for vehicles. Constraints (6g) – (6j) define daily relationships between served stores in two consecutive days, i.e., if node $i$ is visited at day $s-1$, then the related variables to this node at day $s$ are bounded to zero. (6k) – (6l) impose an upper bound on the number of stores that were not visited in the previous days. At last, constraints (6m) reinforce consecutive active days.
@@ 2.4 Upper bounds
where $\ell_{S}$, $\ell_{K}$, and $\ell_{E}$ are decision variables representing the number of needed days, vehicles, and pollsters, respectively. The optimal solution $(\ell_{S}^{*},\ell_{K}^{*},\ell_{E}^{*})$ of (7a)–(7d) provides an upper bound on the objective function (1) of model IVPRM. As a result, this solution determines a tighter value on the cardinality of sets $S$, $K$, and $E$ required to get an aggregated daily plan.
Observe that constraint (7b) distributes a day in two portions. On the one hand, the first term averages the pollstering time and breaks each pollster would take on all days to finish an aggregated daily plan. On the other hand, the second term averages the time required, in a worst-case scenario, to travel by vehicle to leave pollsters and pick them up at each node. The sum of both terms in any feasible solution for IVPRM necessarily has to be below the time horizon $B_{\max}$. Furthermore, constraint (7c) bounds the number of pollsters by vehicle capacity. At last, the optimal solution can be obtained, in the worst case, using exhaustive enumeration in $\mathcal{O}\big{(}|S||K||E|\big{)}$ time.
@@ 2.5 Lower bounds
Likewise, it is possible to determine lower bounds from store visiting times. To do so, first, let $\ell_{e}$ be the lowest integer value that represents the minimum number of pollsters required, in at most $|S|$ days, to distribute the aggregated service time with respect to the maximum shift and mandatory break lengths. This value is obtained by solving the nonlinear program:
@@
Notice that if problem (8) is infeasible, then there are not enough resources to solve IVPRM; thus, providing an infeasibility certificate.
Next, the minimum number of days required to visit all stores, noted $\ell_{s}$, is recovered as the first index $s$ such that $\ell_{e}$ belongs to the set $\Lambda_{E,s}=\{s,s+1,\ldots,s\cdot|E|\}$ for all $s\in\{1,\ldots,|S|\}$. Here $\Lambda_{E,s}$ is the set containing the cumulative number of pollsters that can be active on a day with index $s$.
At last, a bound on the number of vehicles is derived by observing that if a day is active then there must be an active vehicle as well (see (3f)), and that each vehicle can transport at most $Q$ pollsters. Thus the smallest number $\ell_{k}$ that satisfies these constraints solves the following linear program: $\min\big{\{}\ell_{k}:\,\ell_{e}\leq Q\ell_{k}\,\wedge\,\ell_{k}\in\{\ell_{s},\ldots,\ell_{s}|K|\}\big{\}}$.
A solution for these three problems can be obtained with exhaustive enumeration in $\mathcal{O}\big{(}|S||E|+|K|\big{)}$ time. Moreover, while the upper bounds $(\ell_{S},\ell_{K},\ell_{E})$ would change according to the costs in the objective function, the lower bounds $(\ell_{s},\ell_{k},\ell_{e})$ are cost-independent. They are added to IVPRM as the following constraints:
@@
These result in the lower bound $\kappa_{0}\ell_{s}+\kappa_{1}\ell_{k}+\kappa_{2}\ell_{e}$ with respect to the objective of IVPRM.
@@ 2.6 An illustrated example
In order to explain the solution approach for the IVPRP, a small instance is presented in Figure 3 with the following parameters: $n=4$ stores, $|E|=2$ pollsters, $|K|=2$ vehicles with capacity $Q=2$, and a limit of $|S|=2$ days; the break is fixed to one minute within a time frame of $[T_{0},T_{1}]=[20,30]$; the time horizon is set to 30 minutes, the fixed and variable costs are $\kappa_{0}=\$300$, $\kappa_{1}=\$100$, and $\kappa_{2}=\$80$, respectively. The vehicular and pedestrian travel times are shown in Table 2. The diagonal of the pollster travel time matrix corresponds to the service times. Observe that this table does not contain information related to nodes in the set $\{5,6,7,8,9\}$, which are the duplicated nodes explained in the multigraph construction.
@@ 00 $1$ $2$ $3$ $4$ $1$ $2$ $3$ $4$
As portrayed in Figure 3, the optimal solution for this small instance consists of one vehicular route: $0\to 1\to 4\to 8\to 3\to 7\to 6\to 9$ (continuous lines); three pedestrian routes: $1\to 5\to 2\to 6$ for pollster 00, and $4\to 8$ and $3\to 7$ for pollster $1$. Synced pedestrian and vehicular routes form feasible service routes corresponding to the optimal daily plan. Breaks are represented by dashed nodes $2$ and $3$ for pollsters 0 and 1, respectively. The optimal value is accomplished in one day using one vehicle and two pollsters, costing \$$\$560$.
@@ 3.1 Three-phase algorithm
Given that the IVPRP is $\mathcal{NP}$–hard, preliminary tests have shown that finding an optimal solution for instances with even a small number of nodes is not guaranteed in a reasonable time. Thus, a heuristical procedure is proposed in this subsection. The method consists of three phases solved consecutively. In the first phase, the nodes are partitioned into an even number of balanced subsets. Each partition is associated with a part-time daily plan, whether in the morning or afternoon, without breaks. The second phase takes the graph induced by the nodes of each partition as input and solves the IVPRM problem. This input-reduced version of the main problem is notated by (R-IVPRM). The last phase links adequately the solutions found for each partition to obtain feasible daily plans. Other approaches using graph-partitioning techniques have been similarly reported in the literature; see [20, 42].
@@
In order to separate the node set into smaller-size subsets, the following problem is solved: let $G=(C_{-},F)$ be an undirected complete graph with costs on the edges, which coincides with the pedestrian travel times between each pair of stores. Also, let $t:C_{-}\longrightarrow\mathbb{R}^{+}$ a weight function representing service times at stores, and let $k\geq 2$ be a fixed integer number. A $k$-partition of $G$ is a collection of $k$ subgraphs $(C_{1},F(C_{1})),\ldots,(C_{k},F(C_{k}))$ of $G$, where $\{C_{1},\ldots,C_{k}\}$ is a partition of $C_{-}$. Moreover, let $W_{L},W_{U}\in\mathbb{R}^{+}$, $W_{L}\leq W_{U}$, be the lower and upper bounds, respectively, for the total service time of each subgraph which is defined as the sum of the service time of each node of the subgraph. Then, the problem consists in finding a $k$-partition such that $W_{L}\leq\sum_{j\in C_{i}}t_{j}\leq W_{U}$, for all $i\in\{1,\ldots,k\}$, and the total edge cost over all subgraphs is minimized. In order to solve this partitioning problem, the solution approach proposed in [43] is used, where the size constraints are relaxed. Hence, $W_{L}$ and $W_{U}$ are computed as $\mu\frac{n}{k}-\sigma$ and $\mu\frac{n}{k}+\sigma$, respectively, where $\mu$ and $\sigma$ are the sample mean and standard deviation of the service times associated to all the stores.
@@
For any partition obtained in the previous phase, a reduced version of IVPRM is solved to find feasible vehicular and pedestrian routes. The reduced version R-IVPRM consists of the original model setting $|S|=1$, $P=0$, and removing constraints (5a) – (5c). Additionally, simplified versions of the symmetry-breaking inequalities, which do not consider consecutive days, are included, namely (6a), (6b), (6d), and (6e).
A solution of the relaxed LP of IVPRM is used to compute lower bounds on the optimal objective value, and it should be as high as possible. As a way to improve these lower bounds, some valid inequalities are derived below.
@@
Inequality (10a) is valid since the latest arrival time to the depot is greater than the minimum service time needed to collect all data with the maximum number of pollsters, i.e. , $\sum_{i\in C_{-}}{t_{i,i+n}}\big{/}{|E|}\leq B_{\max}$. On the other hand, the total time spent at each store is greater or equal than the total time required to collect data, i.e. , $\sum_{i\in C_{-}}t_{i,i+n}\leq\sum_{i\in C_{-}}(B_{i+n}-B_{i})$. Despite the fact that constraints (4b)–(4e) imply (10b), they are not necessarily true when subcycles appear in fractional solutions of the linear relaxation. Hence, as in a feasible solution it is only possible to visit node $i+n$ after visiting $i$, and $12n+$ 1 is the last visited node, then (10b) hold.
@@
As every partition corresponds to a part-time daily plan, the purpose of this phase is to match two of these subsets in a full-time feasible daily plan. Observe that as each part-time daily plan may contain a different number of active vehicles and pollsters, one can assume that any of them not needed for the afternoon shift can stay at the depot; otherwise, if the workforce is not enough to fulfill the shift, then some vehicles or pollsters are allowed to join the workforce. To avoid inconsistencies, breaks are assigned at the depot between each half-shift, as each day can be split into two half-shifts by setting a new time frame upper bound $\tilde{B}_{\max}=(B_{\max}-P)/2$.
Thus, the following two cases arise for consideration: (a) the required number of vehicles or pollsters differ in at least two sets, and (b) the required number of vehicles and pollsters is the same for all sets in the partition. For the former, the cost-differing sets are linked by matching their costs in descending order by pairs. For the latter, let $\mathcal{K}=(V,E)$ be a complete graph, where each node in $V=\{1,\dots,k\}$ is associated with each feasible solution of R-IVPRM for any subset obtained during the Partitioning procedure. A cost for each edge $\{i,j\}\in E$ is defined as the sum of the arrival time at the depot for the last vehicle at partitions $i$ and $j$. Then, a minimum cost matching in $K$ is found. Notice that if ties arise in (a), these are resolved as a particular case of (b). In either case, a feasible aggregated daily plan is returned.
@@
Feasible daily plans are constructed according to Algorithm 1 which connects the phases explained previously.
@@
while $f$ is false do
     Perform $N_{\max}$ iterations of the partitioning procedure to find a $k$-way
     partition of $V(G)$.
     for $i=1$ to $k$ do
         Find the best feasible solution of R-IVPRM induced by subset $i$,
         in time $T_{\max}$.

### v6

@@ 2.2 Formulation
For the sake of simplicity, the integrated vehicle and pollster routing model (IVPRM) is presented in the following blocks. Table 1 summarizes all the relevant parameters used in the model.
Table 1: Relevant parameters and sets of the proposed model.
Symbol Parameter
$n$ Number of stores
$K$ Set of vehicles
$E$ Set of pollsters
$S$ Set of days in the time horizon
$B_{\max}$ Day length
$P$ Length of the break for pollsters
$Q$ Vehicle capacity
$G\coloneqq(V,A)$ Multi-graph
$V\coloneqq C\cup\{0,2n+1\}$ Node set of $G$
$C\coloneqq C_{-}\cup C_{+}$ Set of nodes associated with stores
$C_{-}\coloneqq\{1,\dots,n\}$ Set of nodes representing the start of service
$C_{+}\coloneqq\{n+1,\dots,2n\}$ Set of nodes representing the end of service
$A\coloneqq A_{S}\cup A_{W}\cup A_{V}$ Arc set of $G$
$A_{S}$ Service arcs connecting each store with its duplicate
$A_{W}$ Walking arcs
$A_{V}$ Vehicle arcs
$t_{a}$ Time spent by a pollster collecting data at a store, for $a\in A_{S}$
$t_{a}$ Pedestrian travel time between stores, for $a\in A_{W}$
$\tau_{a}$ Vehicular travel time between stores, for $a\in A_{V}$
$\kappa_{0}$ Fixed operational cost per working day
$\kappa_{1}$ Fixed vehicle hiring cost
$\kappa_{2}$ Fixed pollster hiring cost
$[T_{0},T_{1}]$ Time window in which pollsters must start their break
$M$ Sufficiently large number
@@
The following sets of binary variables for arc selection are used in the model. First, variable $x_{i,j}^{e,s}$ is equal to one if and only if the arc $(i,j)\in A_{S}\cup A_{W}$ is chosen to be part of some walking path of pollster $e\in E$ on day $s\in S$. Second, variable $y_{i,j}^{k,s}$ is equal to one whenever arc $(i,j)\in A_{V}$ is part of the route of vehicle $k\in K$ on day $s\in S$. Third, $z_{i,j}^{e,s}$ indicates whether pollster $e\in E$ is transported by some vehicle over arc $(i,j)\in A_{V}$ on day $s\in S$. Fourth, $b_{i}^{e,s}$ (correspondingly $f_{i}^{e,s}$) indicates whether a pollster $e\in E$ starts (finishes) a walking path at node $i\in C_{-}$ ($i\in C_{+}$) on day $s\in S$. Fifth, $w_{i}^{e,s}$ indicates whether a pollster $e\in E$ takes a break at node $i\in C_{-}$ on day $s\in S$. Finally, $u_{s}$ determines whether any workload is assigned to pollsters and vehicles on day $s\in S$ for the final aggregated daily plan. Additionally, for each node $i\in C_{-}$, let $B_{i}$ and $B_{i+n}$ represent the arrival and departure times of a pollster at nodes $i$ and $i+n$, respectively.
@@
The objective function aims to minimize the number of operational days and the number of vehicles and pollsters hired. The first component represents the fixed operational daily cost, while the other two components represent the variable costs associated with the number of pollsters and vehicles hired for an operational day.
Although this definition of an objective function is uncommon in Routing Problems, it is justified in this case because the distance walked by a pollster or the travel time spent by a vehicle are irrelevant quantities for this particular application. Instead, the focus is on the fixed daily costs of hiring vehicles and pollsters.
@@
The following set of constraints describes the routes for each pollster $e\in E$ within each day $s\in S$ in the time horizon.
@@
Constraints Equation 2a ensure that each store $i\in C_{-}$ is visited exactly once by a pollster on a given day. Constraints Equation 2b and Equation 2c identify the edges of walking paths; specifically, if $b_{i}^{e,s}=1$ (resp. $f_{i}^{e,s}=1$), then a walking path must begin (resp. end) at node $i\in C_{-}$ (resp. $C_{+}$). Moreover, constraints Equation 2dEquation 2g guarantee that if a node $i$ is the beginning or end of a walking path, then exactly one vehicle must pick up or deliver one pollster at that node. Finally, constraints Equation 2h enforce that each pollster departs from the depot at most once on every active day.
@@
The following constraints define vehicular routes for pollster transportation.
@@
Constraints Equation 3a and Equation 3b ensure that vehicles arrive only at pickup nodes and depart only from delivery nodes. Constraints Equation 3c and Equation 3d are flow conservation constraints. Furthermore, Equation 3e ensure that each vehicle departs from the depot at most once on every active day. Finally, Equation 3f establish the linkage between pollster and vehicular routes.
Time management and shift length
The following constraints are introduced to define time bounds for the duration of a pollster shift as well as the times required to travel and visit each store.
@@
Constraints Equation 4a require that the departure time from a store $i$ be at least as large as the arrival time plus the service time and, if applicable, the time needed for a pause at that store. Note that the sum in the right-most term of these constraints equals one if a pause is taken at node $i$, and zero otherwise. Constraints Equation 4b and Equation 4c ensure that the arrival time at a store within a route accounts for the pedestrian or vehicular travel time from the previously visited store, respectively. If node $i$ immediately follows the depot in any route, then constraints Equation 4d provide a lower bound on the arrival time of vehicles and pollsters. Finally, constraints Equation 4e link the variables $B_{i}$ and $u^{s}$ to bound the latest arrival time of any vehicle at the depot for each day $s$. Here, $M$ is a sufficiently large number such that $M>\max\{t,\tau\}+B_{\max}$.
@@
The following constraints ensure that each pollster on duty takes a break starting within the prescribed time frame on a given day.
@@
Constraints Equation 5a establish that if a break is taken by a pollster at node $i$, it must occur within the time interval $[T_{0},T_{1}]$. Constraints (5b) guarantee that a pollster takes a break only at visited stores, and constraints Equation 5c enforce that each pollster on duty takes a break exactly once per day.
@@ 2.3 Symmetry breaking inequalities
In any feasible aggregated daily plan, a fixed vehicle services a vehicular route; however, the route can be served by any other vehicle in $K$, yielding the same objective value. Similarly, the workload of any pair of pollsters in $E$ can be permuted within a daily plan. Moreover, the same applies to the set of days $S$ concerning the feasible daily plans. As a result, a set of equivalent solutions might arise by relabeling assigned vehicles and pollsters. The following set of constraints aims to avoid these symmetries. This approach can be seen as an application of the well-known technique of symmetry elimination [43, 44, 45].
@@
Constraints Equation 6aEquation 6c ensure an ordered assignment of pollsters, starting with the one labeled as 00, for each day $s\in S$. Similarly, constraints Equation 6dEquation 6f guarantee an ordered assignment for vehicles.
Constraints Equation 6gEquation 6j define daily relationships between served stores across two consecutive days; i.e., if node $i$ is visited on day $s-1$, then the related variables for this node on day $s$ are bounded to zero. Constraints Equation 6kEquation 6l impose an upper bound on the number of stores that were not visited in the previous days. Finally, constraints Equation 6m reinforce the activation of consecutive active days.
@@ 2.4 Upper bounds
where $\ell_{S}$, $\ell_{K}$, and $\ell_{E}$ are decision variables representing the number of needed days, vehicles, and pollsters, respectively. The optimal solution $(\ell_{S}^{*},\ell_{K}^{*},\ell_{E}^{*})$ of Equation 7aEquation 7d provides an upper bound on the objective function Equation 1 of model IVPRM. As a result, this solution determines a tighter value for the cardinality of the sets $S$, $K$, and $E$ required to obtain an aggregated daily plan.
Observe that constraint Equation 7b divides a day into two portions. On the one hand, the first term averages the pollstering time and breaks that each pollster would take across all days to complete an aggregated daily plan. On the other hand, the second term averages the time required, in a worst-case scenario, for vehicles to travel, drop off pollsters, and pick them up at each node. The sum of both terms in any feasible solution for IVPRM must necessarily be below the time horizon $B_{\max}$. Furthermore, constraint Equation 7c limits the number of pollsters according to vehicle capacity. Finally, the optimal solution can be obtained, in the worst case, using exhaustive enumeration in $\mathcal{O}\big{(}|S||K||E|\big{)}$ time.
@@ 2.5 Lower bounds
Likewise, it is possible to determine lower bounds based on store visiting times. To do so, first, let $\ell_{e}$ be the smallest integer value representing the minimum number of pollsters required, in at most $|S|$ days, to distribute the aggregated service time while considering the maximum shift length and mandatory break durations. This value is obtained by solving the following nonlinear program:
@@
Notice that if problem (8) is infeasible, then there are not enough resources to solve IVPRM, thereby providing an infeasibility certificate.
Next, the minimum number of days required to visit all stores, denoted as $\ell_{s}$, is recovered as the first index $s$ such that $\ell_{e}$ belongs to the set $\Lambda_{E,s}=\{s,s+1,\ldots,s\cdot|E|\}$ for all $s\in\{1,\ldots,|S|\}$. Here, $\Lambda_{E,s}$ is the set containing the cumulative number of pollsters that can be active on a day with index $s$.
Finally, a bound on the number of vehicles is derived by noting that if a day is active, then there must also be an active vehicle (see Equation 3f), and each vehicle can transport at most $Q$ pollsters. Thus, the smallest number $\ell_{k}$ that satisfies these constraints solves the following linear program: $\min\big{\{}\ell_{k}:\,\ell_{e}\leq Q\ell_{k}\,\wedge\,\ell_{k}\in\{\ell_{s},\ldots,\ell_{s}|K|\}\big{\}}$.
A solution for these three problems can be obtained through exhaustive enumeration in $\mathcal{O}\big{(}|S||E|+|K|\big{)}$ time. Moreover, while the upper bounds $(\ell_{S},\ell_{K},\ell_{E})$ would change depending on the costs in the objective function, the lower bounds $(\ell_{s},\ell_{k},\ell_{e})$ are cost-independent. These lower bounds are incorporated into IVPRM as the following constraints:
@@
Combining these inequalities yields the lower bound $\kappa_{0}\ell_{s}+\kappa_{1}\ell_{k}+\kappa_{2}\ell_{e}$ with respect to the objective of IVPRM.
@@ 2.6 An illustrated example
To explain the solution approach for the IVPRP, a small instance is presented in Figure 3 with the following parameters: $n=4$ stores, $|E|=2$ pollsters, $|K|=2$ vehicles with a capacity of $Q=2$, and a limit of $|S|=2$ days. The break is fixed at one minute within a time frame of $[T_{0},T_{1}]=[20,30]$, and the time horizon is set to 30 minutes. The fixed and variable costs are $\kappa_{0}=\$300$, $\kappa_{1}=\$100$, and $\kappa_{2}=\$80$, respectively.
@@
The vehicular and pedestrian travel times are shown in Table 2. The diagonal of the pollster travel time matrix corresponds to the service times. Note that this table does not include information related to nodes in the set $\{5,6,7,8,9\}$, which are the duplicated nodes explained in the multigraph construction.
@@ 00 $1$ $2$ $3$ $4$ $1$ $2$ $3$ $4$
As shown in Figure 3, the optimal solution for this small instance consists of one vehicular route: $0\to 1\to 4\to 8\to 3\to 7\to 6\to 9$ (continuous lines), and three pedestrian routes: $1\to 5\to 2\to 6$ for pollster 00, and $4\to 8$ and $3\to 7$ for pollster $1$. Synced pedestrian and vehicular routes form feasible service routes corresponding to the optimal daily plan. Breaks are represented with the dashed nodes $2$ and $3$ for pollsters 00 and $1$, respectively. The optimal value is achieved in one day using one vehicle and two pollsters, with a total cost of \$$\$560$.
@@ 3.1 Three-phase algorithm
Given that the IVPRP is $\mathcal{NP}$–hard, preliminary tests have shown that finding an optimal solution for instances with even a small number of nodes cannot be guaranteed in a reasonable amount of time. Therefore, a heuristic procedure is proposed in this subsection. The method consists of three phases solved consecutively.
In the first phase, the nodes are partitioned into an even number of balanced subsets. Each partition is associated with a part-time daily plan, either in the morning or afternoon, without breaks. The second phase takes the graph induced by the nodes of each partition as input and solves the IVPRM problem. This input-reduced version of the main problem is denoted by (R-IVPRM). The final phase adequately links the solutions found for each partition to construct feasible daily plans.
Other approaches using graph-partitioning techniques have been similarly reported in the literature; see [20, 46].
@@
To divide the node set into smaller subsets, the following problem is solved: let $G=(C_{-},F)$ be an undirected complete graph with edge costs corresponding to the pedestrian travel times between each pair of stores. Additionally, let $t:C_{-}\longrightarrow\mathbb{R}^{+}$ be a weight function representing the service times at stores, and let $k\geq 2$ be a fixed integer. A $k$-partition of $G$ is a collection of $k$ subgraphs $(C_{1},F(C_{1})),\ldots,(C_{k},F(C_{k}))$ of $G$, where $\{C_{1},\ldots,C_{k}\}$ forms a partition of $C_{-}$. Moreover, let $W_{L},W_{U}\in\mathbb{R}^{+}$, with $W_{L}\leq W_{U}$, be the lower and upper bounds for the total service time of each subgraph, defined as the sum of the service times of all nodes within the subgraph. The problem consists of finding a $k$-partition such that $W_{L}\leq\sum_{j\in C_{i}}t_{j}\leq W_{U}$ for all $i\in\{1,\ldots,k\}$, while minimizing the total edge cost over all subgraphs.
To solve this partitioning problem, the approach proposed in [47] is used, where the size constraints are relaxed. Hence, $W_{L}$ and $W_{U}$ are computed as $\nicefrac{{\mu n}}{{k}}-\sigma$ and $\frac{\mu n}{k}+\sigma$, respectively, where $\mu$ and $\sigma$ are the sample mean and standard deviation of the service times associated with all the stores.
@@
For any partition obtained in the previous phase, a reduced version of IVPRM is solved to find feasible vehicular and pedestrian routes. The reduced version, denoted as R-IVPRM, consists of the original model with $|S|=1$, $P=0$, and the removal of constraints Equation 5aEquation 5c. Additionally, simplified versions of the symmetry-breaking inequalities, which do not consider consecutive days, are included; specifically, Equation 6aEquation 6b, and Equation 6dEquation 6e.
A solution of the relaxed LP of IVPRM is used to compute lower bounds on the optimal objective value, and this value should be as tight as possible. To improve these lower bounds, some valid inequalities are derived below.
@@
Inequality Equation 10a is valid because the latest arrival time to the depot is greater than the minimum service time required to collect all data using the maximum number of pollsters; i.e., $\sum_{i\in C_{-}}{t_{i,i+n}}\big{/}{|E|}\leq B_{\max}$. On the other hand, the total time spent at each store is greater than or equal to the total time required to collect data; i.e., $\sum_{i\in C_{-}}t_{i,i+n}\leq\sum_{i\in C_{-}}(B_{i+n}-B_{i})$.
Although constraints Equation 4bEquation 4e imply Equation 10b, they are not necessarily valid when subcycles appear in fractional solutions of the linear relaxation. Hence, as in a feasible solution, it is only possible to visit node $i+n$ after visiting $i$, and $12n+$ 1 must be the last visited node. Therefore, Equation 10b holds.
@@
Since each partition corresponds to a part-time daily plan, the purpose of this phase is to match two of these subsets into a full-time feasible daily plan. Note that, as each part-time daily plan may involve a different number of active vehicles and pollsters, it can be assumed that any vehicles or pollsters not needed for the afternoon shift can remain at the depot. Conversely, if the workforce is insufficient to fulfill the shift, additional vehicles or pollsters may join the workforce. To avoid inconsistencies, breaks are assigned at the depot between each half-shift, as each day can be split into two half-shifts by setting a new time frame upper bound $\tilde{B}_{\max}=(B_{\max}-P)/2$.
Thus, the following two cases arise for consideration: (a) the required number of vehicles or pollsters differs between at least two sets, and (b) the required number of vehicles and pollsters is the same for all sets in the partition. For the former, the cost-differing sets are linked by matching their costs in descending order by pairs. For the latter, let $\mathcal{K}=(V,E)$ be a complete graph, where each node in $V=\{1,\dots,k\}$ is associated with a feasible solution of R-IVPRM for any subset obtained during the Partitioning procedure. A cost for each edge $\{i,j\}\in E$ is defined as the sum of the arrival times at the depot for the last vehicle in partitions $i$ and $j$. Then, a minimum-cost matching in $\mathcal{K}$ is found. Note that if ties arise in case (a), they are resolved as a special instance of case (b). In either case, a feasible aggregated daily plan is returned.
@@
Feasible daily plans are constructed according to Algorithm 1, which connects the phases explained previously.
@@
while $f=\texttt{false}$ do
     Perform $N_{\max}$ iterations of the partitioning procedure to find a $k$-way partition
     of $V(G)$.
     for $i\in\{1,\ldots,k\}$ do
         Find the best feasible solution of R-IVPRM induced by subset $i$, within the
         time limit $T_{\max}$.